Consider the following zero-coupon yields on default free securities:
Maturity (years) Zero-Coupon YTM 1 2. 3 5.80% 5.50% 5.20% 5.00% 4.80%

A 4 year default free security with a face value of $1000 and an annual coupon rate of 5.25% will trade:

Group of answer choices

at a premium.

at par.

at a discount.

There is insufficient information provided to answer this question.
Answer #1

a) face value = F= 1000

coupon rate = 5.25%

coupon payment = C = 52.5

time period = n = 4 years

YTM = 5%

since coupon rate > YTM, we can directly say that bond is trading at premium. But I can also show it with the proof by calculating price below

using bond price formula,

p = C1/(1+YTM)1 + C2/(1+YTM)2 + C3/(1+YTM)3 + (C4+F)/(1+YTM)4

= 52.5/(1.05)1 + 52.5/(1.05)2 + 52.5/(1.05)3 + 1052.5/(1.05)4

= 50 + 47.62 + 45.35 + 865.89

= 1008.86

since price > face value, the bond is trading at premium